# Fermion masses and mixings and charged lepton flavor violation in a 3-3-1 model with inverse seesaw

### v1

@@ V Higgs and gauge bosons
The model consists of $8$ electroweak gauge bosons $W^{a}_{\mu}$ of the $SU(3)_{L}$ gauge group and the $X_{\mu}$ of the $U(1)_{X}$. The corresponding covariant derivative is:
@@
$\frac{g_{X}}{g}=\frac{3\sqrt{2}t_{W}}{\sqrt{3-t_{W}^{2}}}$ (76)
@@
After symmetry breaking, the model contains SM gauge bosons, including the massless photon $A_{\mu}$, as well as the neutral and singly charged gauge bosons $Z$ and $W^{\pm}$.
$W^{\pm}_{\mu}\equiv\frac{W^{1}_{\mu}\mp iW^{2}_{\mu}}{\sqrt{2}},\quad m_{W}=\frac{gv_{\rho}}{2},$ (77)
and in the limit $v^{2}_{\rho}/v^{2}_{\chi}\sim O(\lambda^{6})\simeq 0$, we have $m^{2}_{Z}\simeq m^{2}_{W}/c_{W}^{2}$, which is consistent with the SM prediction. There are five heavy gauge bosons, namely the neutral heavy gauge boson $Z^{\prime}$ with
$m^{2}_{Z^{\prime}}\simeq\frac{g^{2}v_{\chi}^{2}}{3-t_{W}^{2}},$ (78)
and
$X^{0}_{\mu}\equiv\frac{W^{4}_{\mu}-iW^{5}_{\mu}}{\sqrt{2}},\quad Y^{\pm}_{\mu}\equiv\frac{W^{6}_{\mu}\mp iW^{7}_{\mu}}{\sqrt{2}},m_{X}=\frac{gv_{\chi}}{2},\quad m^{2}_{Y}=m^{2}_{X}+m^{2}_{W}.$ (79)
The $Z^{\prime}$ relate to the original basis $(X_{\mu},W^{3}_{\mu},W^{8}_{\mu})$ through the following transformation:
$\begin{pmatrix}X_{\mu}\\ W^{3}_{\mu}\\ W^{8}_{\mu}\end{pmatrix}=C_{zz^{\prime}}\begin{pmatrix}A_{\mu}\\ Z_{\mu}\\ Z^{\prime}_{\mu}\end{pmatrix},\quad C_{zz^{\prime}}=\left(\begin{array}[]{ccc}c_{W}s_{331}&c_{331}s_{z}-c_{z}s_{331}s_{W}&c_{331}c_{z}+s_{331}s_{W}s_{z}\\ s_{W}&c_{W}c_{z}&-c_{W}s_{z}\\ c_{331}c_{W}&-c_{331}c_{z}s_{W}-s_{331}s_{z}&c_{331}s_{W}s_{z}-c_{z}s_{331}\\ \end{array}\right),$ (83)
@@
The details of the Higgs potential, see Appendix B. The model contains a singly charged Higgs boson, and a SM-like Higgs boson found by LHC.
VI Lepton flavor violation decays $e_{j}\rightarrow e_{i}$
@@
Given that the one-loop contributions from neutral Higgs bosons, $Z$ and $Z^{\prime}$ bosons involve couplings only to the SM charged lepton, we will be ignored them in this work. Therefore, we will focus solely on the couplings of charged leptons to both heavy neutral leptons and charged gauged bosons, which are relevant for one-loop contributions to LFV decays. The calculation will be done in the unitary gauge.
The left-handed charged leptons are transformed from the flavor basis $l_{iL,R}$ to the physical basis $e_{iL,R}$ using the matrix $R_{\ell L}$ from Eq. (27) $l_{iL}=R_{\ell L,ij}e_{jL}$ with $i,j=1,2,3$. The neutrino mass matrix $M_{\nu}$ in Eq. (40) is diagonalized by the unitary matrix $U_{\nu}$:
@@
Processes like $\mu\rightarrow e\gamma$ are crucial for probing beyond SM physics, including 3-3-1 models. Based on the developments presented in [125, 125, 126], we explore the LFV processes arising from non-universal couplings of the standard charged leptons to the gauge bosons. In our model, LFV is directly linked to the masses and interactions of exotic neutrinos and their associated gauge bosons. Experimental limits, particularly MEG ones, impose stringent constraints on the parameter space of 3-3-1 models. Future experiments, conversion in aluminum and titanium nuclei, will further refine these constraints and provide insights into the scale of new physics.
Figure 3: The Branching ratio for $\mu\rightarrow e\gamma$ as a function of the mass of the gauge boson Y. The upper shaded region corresponds to an area excluded by the MEG [127] and the lower shaded region corresponds to the expected sensitivities of the next generation of experiments using aluminium as targets [128].
The analytic expression for the branching ratio (Br) for the decay of $e_{i}\rightarrow e_{j}\gamma$ ($m_{e_{i}}>m_{e_{j}}$) is given by [125]:
@@
where $G_{F}=g^{2}/(4\sqrt{2}m_{W}^{2})$ is the Fermi constant, and $\Gamma_{ij}$ represents the one-loop contribution from virtual charged gauge bosons and Majorana neutrinos in the loop diagrams. The contribution can be expressed as: $\Gamma_{ij}=\Gamma^{W}_{ij}+\Gamma^{Y}_{ij}$, where:
@@
$\Gamma^{Y}_{ij}$ $=-\frac{eg^{2}}{32\pi^{2}m_{Y}^{2}}\sum_{k=1}^{9}\sum_{a,b=1}^{3}(R_{lL}^{*})_{ib}(U^{*}_{\nu})_{(b+3)k}(R_{lL})_{ja}(U_{\nu})_{(a+3)k}\mathcal{F}\left(\frac{m^{2}_{n_{k}}}{m_{Y}^{2}}\right).$
The loop function $\mathcal{F}(z)$ is defined as:
$\mathcal{F}(z)=2(z+2)\mathcal{I}_{3}(z)-2(2z-1)\mathcal{I}_{2}(z)+2z\mathcal{I}_{1}(z)+1,$ (92)
@@
Figure 3 displays the model predictions for the branching ratio $\text{Br}(\mu\rightarrow e\gamma)$ as a function of the exotic gauge boson mass $m_{Y}$. The color-coded points represent different exotic neutrino mass values: green for $1$ TeV, gray for $2$ TeV, and blue for $5$ TeV. The yellow shaded region, constrained by the MEG experiment’s limit of $\text{Br}(\mu\rightarrow e\gamma)<4.2\times 10^{-13}$ [127], excludes certain parameter space regions. The cyan shaded region indicates the projected sensitivity of future experiments like $\mu-e$ conversion in titanium and aluminum nuclei, with a sensitivity of $\text{CR}(\mu^{-}\mathrm{Al}\rightarrow e^{-}\mathrm{Al})\lesssim 10^{-17}$ [128]. Numerically, the relevant couplings are given by $x_{\rho}=1.99$ and $y_{1\chi}^{\left(L\right)}=1.92$, which directly affect the magnitude of the contributions from gauge bosons and exotic neutrinos to LFV processes, since they control the active-sterile neutrino mixing angles given by Eq. (88), which determine the couplings of the neutrinos with SM charged leptons and charged gauge bosons, then giving rise to charged lepton flavor violating processes.
@@
In Figure 4, the correlations between the branching ratios $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$, as well as between $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu\text{Al}\to e\text{Al})$, are shown. In panel (a), the relationship between $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$ is illustrated through a color gradient that indicates the values of $\text{Br}(\tau\to\mu\gamma)$. This panel provides insight into how lepton flavor violation in the muon and tau decays correlates with the respective branching ratios of these processes. In panel (b), the correlation between $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$ is presented, with a color gradient reflecting the values of the gauge boson mass $m_{Y}$. This offers a direct indication of the dependence of the conversion rates on the mass of the new gauge boson in our model. The yellow shaded region in both panels marks the exclusion zone from the experimental limits set by the MEG experiment [127].
@@
Finally, in Table 5, we present numerical cases for the heavy neutrino masses $m_{n_{4}}$ to $m_{n_{9}}$, also includes the mass of the gauge boson $m_{Y}$ in TeV, the corresponding branching ratios for processes $\mu\to e\gamma$, $\tau\to e\gamma$ and $\tau\to\mu\gamma$ and the estimated conversion rate $\mathrm{CR}(\mu^{-}\mathrm{Al}\to e^{-}\mathrm{Al})$. These values provide insight into the correlations between the LFV processes and the masses of the heavy neutrinos and gauge bosons within the context of our model.

### v2

@@ V Higgs and gauge bosons
The model consists of $8$ electroweak gauge bosons $W^{a}_{\mu}$ of the $SU(3)_{L}$ gauge group and $X_{\mu}$ of $U(1)_{X}$. The corresponding covariant derivative is:
@@
$\frac{g_{X}}{g}=\frac{3\sqrt{2}t_{W}}{\sqrt{3-t_{W}^{2}}},$ (80)
@@
After symmetry breaking, the model contains SM gauge bosons, including the massless photon $A_{\mu}$with their field given by
$A_{\mu}=s_{W}W_{3\mu}-\frac{s_{W}}{\sqrt{3}}W_{8\mu}+\frac{\sqrt{1+2c_{2W}}}{\sqrt{3}}X_{\mu},$ (81)
as well as two neutral weak gauge bosons
$Z_{\mu}=c_{W}W_{3\mu}-t_{W}\left(-\frac{s_{W}}{\sqrt{3}}W_{8\mu}+\frac{\sqrt{1+2c_{2W}}}{\sqrt{3}}X_{\mu}\right),\hskip 14.22636ptZ^{\prime}_{\mu}=\frac{\sqrt{1+2c_{2W}}}{\sqrt{3}s_{W}}W_{8\mu}+\frac{t_{W}}{\sqrt{3}}X_{\mu},$ (82)
with $c_{W}=\cos\theta_{W},s_{W}=\sin\theta_{W}$. In the former expressions, the neutral currents of $Z_{\mu}$ coincide with the weak neutral current of the SM. However, in the model considered, the physical fields $Z_{1\mu},Z_{2\mu}$ are defined
$Z_{1\mu}=c_{z}Z_{\mu}-s_{z}Z^{\prime}_{\mu},\hskip 14.22636ptZ_{2\mu}=s_{z}Z_{\mu}+c_{z}Z^{\prime}_{\mu}$ (83)
@@
In the limit $\frac{v^{2}_{\rho}}{v^{2}_{\chi}}\sim O(\lambda^{6})\simeq 0$, we have
$m^{2}_{Z_{1}}\simeq\frac{m^{2}_{W}}{c_{W}^{2}},\hskip 14.22636ptm^{2}_{Z_{2}}\simeq\frac{g^{2}v_{\chi}^{2}}{3-t_{W}^{2}}.$ (85)
The single-charged gauge bosons $Z$ and $W^{\pm}$.
$W^{\pm}_{\mu}\equiv\frac{W^{1}_{\mu}\mp iW^{2}_{\mu}}{\sqrt{2}},\quad m_{W}=\frac{gv_{\rho}}{2}.$ (86)
There are four new heavy-complex gauge bosons
$X^{0}_{\mu}\equiv\frac{W^{4}_{\mu}-iW^{5}_{\mu}}{\sqrt{2}},\hskip 28.45274ptY^{\pm}_{\mu}\equiv\frac{W^{6}_{\mu}\mp iW^{7}_{\mu}}{\sqrt{2}},$ (87)
with their masses
$m_{X}=\frac{gv_{\chi}}{2},\hskip 28.45274ptm^{2}_{Y}=m^{2}_{X}+m^{2}_{W}.$ (88)
For user convenience, we provide the basis transformation matrix of the neutral gauge bosons, defined as
$\begin{pmatrix}X_{\mu}\\ W^{3}_{\mu}\\ W^{8}_{\mu}\end{pmatrix}=C_{zz^{\prime}}\begin{pmatrix}A_{\mu}\\ Z_{1\mu}\\ Z_{2\mu}\end{pmatrix},\quad C_{zz^{\prime}}=\left(\begin{array}[]{ccc}c_{W}s_{331}&c_{331}s_{z}-c_{z}s_{331}s_{W}&c_{331}c_{z}+s_{331}s_{W}s_{z}\\ s_{W}&c_{W}c_{z}&-c_{W}s_{z}\\ c_{331}c_{W}&-c_{331}c_{z}s_{W}-s_{331}s_{z}&c_{331}s_{W}s_{z}-c_{z}s_{331}\\ \end{array}\right),$ (92)
@@
For the details of the Higgs potential, see Appendix B. The model contains a singly charged Higgs boson and a SM-like Higgs boson found by LHC.
VI FCNCs
Since the quark generations are non-universal under the $SU(3)_{L}\times U(1)_{X}$ gauge symmetry, the corresponding tree-level flavor-changing neutral currents (FCNCs) must also be non-universal. Indeed, considering the interaction of neutral gauge bosons with fermions, we have:
$\mathcal{L}_{N.C}=-g\bar{f}\gamma^{\mu}\left\{T_{3}W_{3\mu}+T_{8}A_{8\mu}+\frac{t}{\sqrt{6}}\left(Q-T_{3}-\beta T_{8}\right)B_{\mu}\right\}f,$ (93)
where $f$ runs over all fermion multiplets. The FCNCs associate only with $T_{8}$ for ordinary quarks, the revenant interactions are
$\mathcal{L}_{N.C}$ $\supset$ $-g\bar{q}_{aL}\gamma^{\mu}T_{8}q_{aL}\left(A_{8\mu}-\frac{\beta}{\sqrt{6}}tB_{\mu}\right)=-\frac{g\sqrt{3}}{\sqrt{3-t_{W}^{2}}}\bar{q}_{aL}\gamma^{\mu}T_{8}q_{aL}Z_{\mu}^{\prime}$ (94)
$=$ $-\frac{g\sqrt{3}}{\sqrt{3-t_{W}^{2}}}\bar{q}_{L}\gamma^{\mu}T_{q}q_{L}\left(-s_{z}Z_{1\mu}+c_{z}Z_{2\mu}\right),$
with $T_{q}=\frac{1}{2\sqrt{3}}\text{Diag}\left(1,-1,-1\right)$ for gauge states $q=u_{L}=\left(u_{1L},u_{2L},u_{3L}\right)^{T}$ or $q=d_{L}=\left(d_{1L},d_{2L},d_{3L}\right)^{T}$. The gauge states are related to the mass eigenstates by $u_{L,(R)}=V_{u_{L}(u_{R})}u^{\prime}_{L,(R)}$ and $d_{L,(R)}=V_{d_{L}(u_{R})}d^{\prime}_{L,(R)}$. The CKM matrix is given by $V_{\text{CKM}}=V_{u_{L}}^{\dagger}V_{d_{L}}$. Due to the specific form of the quark mass matrix presented in Eq. (26), the mixing matrix $V_{d_{L}}$ is simply the identity matrix, specifically $V_{d_{L}}=\text{Diag}\left(1,1,1\right)$. This indicates that $V_{u_{L}}^{\dagger}$ is equivalent to the CKM matrix, which means $V_{u_{L}}^{\dagger}=V_{\text{CKM}}$. Consequently, transitioning to the mass eigenstates results in tree-level FCNCs only happen in $u$-quark sector
$\mathcal{L}_{\text{FCNC}}$ $=$ $-\frac{g}{\sqrt{3-t_{W}^{2}}}\left\{\left(V^{T}_{\text{CKM}}\right)_{1i}\left(V_{\text{CKM}}^{\dagger}\right)_{1j}\right\}\bar{u}^{\prime}_{iL}\gamma^{\mu}u_{jL}\left(-s_{z}Z_{1\mu}+c_{z}Z_{2\mu}\right)$ (95)
$=$ $\vartheta_{ij}\bar{u}^{\prime}_{iL}\gamma^{\mu}u_{jL}\left(-s_{z}Z_{1\mu}+c_{z}Z_{2\mu}\right),$
with
$\vartheta_{ij}=-\frac{g}{\sqrt{3-t_{W}^{2}}}\left\{\left(V^{T}_{\text{CKM}}\right)_{1i}\left(V_{\text{CKM}}^{\dagger}\right)_{1j}\right\}.$ (96)
Additional, the model also contains the scalar FCNCs associated to neutral scalars content in the triplet scalar Higgs. Two physical states carrying both even $Z_{2}^{L}$ and CP charges, $(h_{1},h_{2})$, relates to two scalar components $(\xi_{\rho},\xi_{\chi})$ by unitary rotation matrix as follows
$\left(\xi_{\rho},\xi_{\chi},\right)^{T}=\mathcal{R}\left(h_{1},h_{2}\right)^{T}.$ (97)
There is a tiny mixing between the two components $\xi_{\rho}$ and $\xi_{\chi}$ because $\xi_{\rho}$ is a doublet of the $SU(2)_{L}$ group, while $\xi_{\chi}$ is a singlet of this group. This implies that the matrix elements $\mathcal{R}_{11}\simeq\mathcal{R}_{22}\sim 1$ and $\mathcal{R}_{12}\simeq\mathcal{R}_{21}\sim\frac{v_{\rho}}{v_{\chi}}\sim\lambda^{2}$. After changing to the physical states, we obtain the scalar FCNCs associated with the CP-even scalar Higgs:
$\mathcal{L}_{\text{FCNC}}^{\text{scalar}}=\Gamma_{\alpha\alpha^{\prime}}^{h_{i}}\bar{u}_{\alpha L}u_{\alpha^{\prime}R}h_{i}+h.c.$ (98)
with
$\Gamma_{\alpha\alpha^{\prime}}^{h_{i}}=\frac{m_{u_{\alpha}^{\prime}}}{v_{\rho}}\left(\mathcal{R}_{1i}+\mathcal{R}_{2i}\lambda^{2}-1\right)\left(V_{\text{CKM}}^{T}\right)_{1\alpha}\left(V_{\text{CKM}}^{\dagger}\right)_{1\alpha^{\prime}}\simeq\frac{m_{u_{\alpha}^{\prime}}}{v_{\rho}}\lambda^{4}\left(V_{\text{CKM}}^{T}\right)_{1\alpha}\left(V_{\text{CKM}}^{\dagger}\right)_{1\alpha^{\prime}}.$ (99)
The coefficients given by Eqs. (96) and (99) show that $\vartheta_{ij}\gg\Gamma_{ij}$. The $\tan 2\theta_{z}\simeq\lambda^{4}$. These mean that the FCNCs associated with the neutral scalars and $Z_{1\mu}$ are suppressed compared to those of the new neutral gauge bosons $Z_{2\mu}$. Hence, we may omit the FCNC contributions related to the neutral Higgs bosons and $Z_{1\mu}$in the following studies.
We want to emphasize that, unlike previous studies, our model predicts the existence of tree-level FCNCs primarily associated with the $u$-quark. This feature is shown in Figure 6, where the Feynman diagram depicts FCNCs in the up-quark sector, mediated by the neutral bosons $Z_{\mu}$ and $Z^{\prime}_{\mu}$. Consequently, our model is not tightly constrained by B-meson oscillations but only by D-meson oscillations. At tree level, the effective Lagrangian for the D-meson oscillation process is as follows:
$\mathcal{L}_{Z^{\prime}\text{eff}}^{D_{0}-\bar{D}_{0}}=\mathcal{G}^{\prime}\frac{m_{Z_{1}}^{2}}{m_{Z_{2}}^{2}}\left|\left(V_{\text{CKM}}\right)_{11}\left(V_{\text{CKM}}^{*}\right)_{21}\right|^{2}\left|\bar{u}^{\prime}_{1L}\gamma^{\mu}u_{2L}^{\prime}\right|^{2},$ (100)
where $\mathcal{G}^{\prime}=\frac{4\sqrt{2}G_{F}}{1+4c_{W}^{2}}$, with $G_{F}$ being the Fermi constant. This effective Lagrangian contributes to the mass splittings $\Delta m_{D}$ between neutral mesons $D_{0}$ and $\bar{D}_{0}$ as follows:
$\Delta m_{D}=\mathcal{G}^{\prime}\frac{m_{Z_{1}}^{2}}{m_{Z_{2}}^{2}}\left|\left(V_{\text{CKM}}\right)_{11}\left(V_{\text{CKM}}^{*}\right)_{21}\right|^{2}f_{D}^{2}B_{D}\eta_{D}m_{D}.$ (101)
Values for the bag parameter $B_{D}$, decay constant $f_{D}$, QCD correction factor $\eta_{D}$, and the D-meson mass are taken from [134]:
$\sqrt{B_{D}}f_{D}=187\text{MeV},\hskip 28.45274pt\eta_{D}=0.57,\hskip 28.45274ptm_{D}=(1865\pm 0.0005)\text{MeV}.$ (102)
The SM predicted value for $\Delta m_{D}$ is
$\left(\Delta m_{D}\right)_{\text{SM}}=10^{-14}\text{MeV},$ (103)
while the experimental value is
$\left(\Delta m_{D}\right)_{\text{exp}}=\left(6.25316^{+2.69873}_{-2.8962}\right)\times 10^{-12}\text{MeV}.$ (104)
This means that the main contribution to the D-meson mass difference comes from the NP. From the experimental bound, we obtain the lower bound on the $Z_{2}$-boson mass: $m_{Z_{2}}>173\text{TeV}$.
We would like to point out that if we organize the third quark family transform differently from the first two families under the symmetry group $SU(3)_{c}\times U(1)_{L}$, the NP contribution to the D-meson mass difference becomes proportional to $\left|\left(V_{\text{CKM}}\right)_{13}\left(V_{\text{CKM}}^{*}\right)_{32}\right|^{2}$ instead of being proportional to $\left|\left(V_{\text{CKM}}\right)_{11}\left(V_{\text{CKM}}^{*}\right)_{21}\right|^{2}$, as considered in the version considered. In this scenario, the constraint on the oscillation of $D_{0}$ and $\bar{D}_{0}$ results in a lower bound on the mass of the new neutral boson, which is approximately a few TeV.
Since the three lepton families transform identically under $SU(3)_{c}\times U(1)_{L}$, lepton FCNCs are forbidden at the tree level but occur at the loop level, as demonstrated in [135]. The model allows for lepton flavor violation (LFV) via one-loop penguin diagrams in meson decays, particularly in D-meson decays. The general four-fermion Hamiltonian describing the interaction for the process $u^{\prime}_{1}\to u^{\prime}_{2}l_{\alpha}^{+}l_{\beta}^{-}$ can be expressed as follows:
$\mathcal{H}_{\text{eff}}=\frac{4G_{\text{F}}}{\sqrt{2}}\left(C_{9}O_{9}+C_{10}O_{10}\right),$ (105)
where
$C_{9}$ $\simeq$ $\frac{2}{1+4c_{W}^{2}}\frac{4\pi^{2}}{e^{2}}\left(\frac{m_{Z_{1}}^{2}}{m_{Z_{2}}^{2}}\right)\left(V_{\text{CKM}}\right)_{11}\left(V_{\text{CKM}}^{*}\right)_{12}\left(\bar{a}_{l}+\bar{a}_{r}\right),$
$C_{10}$ $\simeq$ $\frac{2}{1+4c_{W}^{2}}\frac{4\pi^{2}}{e^{2}}\left(\frac{m_{Z_{1}}^{2}}{m_{Z_{2}}^{2}}\right)\left(V_{\text{CKM}}\right)_{11}\left(V_{\text{CKM}}^{*}\right)_{12}\left(\bar{a}_{l}-\bar{a}_{r}\right).$ (106)
In our case, the expressions for $\bar{a}_{l(r)}$ are determined similarly to those for $a_{l(r)}$ given in [135]. In that work, they predicted $\text{Br}\left(Z\to\mu e\right)\sim 10\times 10^{-12}$, which resulted in coefficient values of $a_{l(r)}\simeq 10^{-3}$. Combining this with the limit $m_{Z_{2}}>173\text{TeV}$, we can roughly estimate the Wilson coefficients $C_{10},C_{9}\sim 10^{-10}$. Due to these small values, we predict suppressed values for the branching ratios of charged lepton flavor violation in D-meson decays, well below the experimental upper bound reported in [136].
Due to the light-heavy neutrino mixing matrix elements, the model also yields the box diagrams for the transition $q_{i}\to q_{j}l_{\alpha}^{+}l_{\beta}^{-}$. The internal lines can be either a light neutrino or heavy neutrino, and either the SM $W_{\mu}^{\pm}$ or new charged gauge bosons $Y_{\mu}^{\pm}$. However, the contributions to the Wilson coefficients from the box diagrams are very small[137, 138]. Hence, we do not consider them in our work.
Figure 6: Feynman diagram representing flavor-changing neutral currents (FCNC) in the up-quark sector. The vertex involves the neutral bosons $Z_{\mu}$ and $Z^{\prime}_{\mu}$.
VII Lepton flavor violating decays $e_{j}\rightarrow e_{i}$
@@
Given that the one-loop contributions from neutral Higgs bosons, $Z$ and $Z^{\prime}$ bosons involve couplings only to the SM charged lepton, we will be ignored them in this work. Therefore, we will focus solely on the couplings of charged leptons to both heavy neutral leptons and charged gauged bosons, which are relevant for one-loop contributions to LFV decays. The calculation will be performed on the unitary gauge.
Left-handed charged leptons are transformed from the flavor basis $l_{iL,R}$ to the physical basis $e_{iL,R}$ using the matrix $R_{\ell L}$ from Eq. (30) $l_{iL}=R_{\ell L,ij}e_{jL}$ with $i,j=1,2,3$. The neutrino mass matrix $M_{\nu}$ in Eq. (43) is diagonalized by the unitary matrix $U_{\nu}$:
@@
Processes like $\mu\rightarrow e\gamma$ are crucial for probing beyond SM physics, including 3-3-1 models. Based on the developments presented in [140, 140, 141]., we explore the LFV processes arising from non-universal couplings of the standard charged leptons to the gauge bosons. In our model, LFV is directly linked to the masses and interactions of exotic neutrinos and their associated gauge bosons. Experimental limits, particularly MEG ones, impose stringent constraints on the parameter space of 3-3-1 models. Future experiments, conversion in aluminum and titanium nuclei, will further refine these constraints and provide insights into the scale of new physics.
Figure 7: Branching ratio for $\mu\rightarrow e\gamma$ as a function of the mass of the gauge boson $Y$ (panel a) and the charged scalar $\rho_{3}^{\pm}$ (panel b). The upper shaded region corresponds to an area excluded by the MEG II experiment [142] and the lower shaded region corresponds to the expected sensitivities of the next generation of experiments using aluminium as targets [143].
The analytic expression for the branching ratio (Br) for the decay of $e_{i}\rightarrow e_{j}\gamma$ ($m_{e_{i}}>m_{e_{j}}$) is given by [140, 144, 145, 146, 147]:
@@
where $\Gamma_{ij}$ represents the one-loop contribution from virtual charged gauge bosons and Majorana neutrinos in the loop diagrams. The contribution can be expressed as: $\Gamma_{ij}=\Gamma^{W}_{ij}+\Gamma^{Y}_{ij}+\Gamma^{\rho_{3}^{\pm}}_{ij}$, where:
@@
$\Gamma^{Y}_{ij}$ $=-\frac{eg^{2}}{32\pi^{2}m_{Y}^{2}}\sum_{k=1}^{9}\sum_{a,b=1}^{3}(R_{lL}^{*})_{ib}(U^{*}_{\nu})_{(b+3)k}(R_{lL})_{ja}(U_{\nu})_{(a+3)k}\mathcal{F}\left(\frac{m^{2}_{n_{k}}}{m_{Y}^{2}}\right),$
$\Gamma^{\rho_{3}^{\pm}}_{ij}$ $=\frac{e^{2}}{2}\sum_{k=1}^{9}\frac{z_{kj}^{*}\,z_{ki}}{(4\pi)^{2}}\frac{1}{m_{\rho_{3}^{\pm}}^{2}}\mathcal{G}\left(\frac{m^{2}_{n_{k}}}{m_{\rho_{3}^{\pm}}}\right)$
@@
$z_{ki}=\sum_{r=1}^{3}\sum_{s=1}^{3}y_{\rho_{3}^{-}}^{rs}\left(R_{lL}^{\dagger}\right)_{ir}\left(\tilde{U}_{\nu}\right)_{sk},$ (114)
and the loop functions $\mathcal{F}(z)$ and $\mathcal{G}(z)$ are defined as:
$\mathcal{F}(z)$ $=$ $2(z+2)\mathcal{I}_{3}(z)-2(2z-1)\mathcal{I}_{2}(z)+2z\mathcal{I}_{1}(z)+1,$
$\mathcal{G}(z)$ $=$ $\frac{1-6z+3z^{2}+2z^{3}-6z^{2}\log z}{6(1-z)^{4}}\,,$ (115)
@@
Figure 7 displays the model predictions for the branching ratio $\text{Br}(\mu\rightarrow e\gamma)$ as a function of the exotic gauge boson mass $m_{Y}$ and the charged scalar mass $m_{\rho_{3}^{\pm}}$. In panel (a) the color-coded points represent different exotic neutrino mass values: green for $1$ TeV, gray for $2$ TeV, and blue for $5$ TeV, whereas in panel (b) the color-coded point represent different mass ranges of the exotic gauge boson $Y$. The yellow shaded region, constrained by the MEG II experiment’s limit of $\text{Br}(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ [142], excludes certain parameter space regions. The cyan shaded region indicates the projected sensitivity of future experiments like $\mu-e$ conversion in titanium and aluminum nuclei, with a sensitivity of $\text{CR}(\mu^{-}\mathrm{Al}\rightarrow e^{-}\mathrm{Al})\lesssim 10^{-17}$ [143]. Numerically, the Yukawa couplings $y_{\rho_{3}^{-}}$ arising from the Dirac block fluctuate within the range $0.01$ to $0.1$, and the relevant couplings are given by $x_{\rho}=1.99$ and $y_{1\chi}^{\left(L\right)}=1.92$, which directly affect the magnitude of the contributions from gauge bosons and exotic neutrinos to LFV processes, since they control the active-sterile neutrino mixing angles given by Eq. (110), which determine the couplings of the neutrinos with SM charged leptons and charged gauge bosons, then giving rise to charged lepton flavor violating processes.
@@
In Figure 8, the correlations between the branching ratios $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$, as well as between $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu\text{Al}\to e\text{Al})$, are shown. In panel (a), the relationship between $\text{Br}(\mu\to e\gamma)$ and $\text{Br}(\tau\to e\gamma)$ is illustrated through a color gradient that indicates the values of $\text{Br}(\tau\to\mu\gamma)$. This panel provides insight into how lepton flavor violation in the muon and tau decays correlates with the respective branching ratios of these processes. In panel (b), the correlation between $\text{Br}(\tau\to\mu\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$ is presented, with a color gradient reflecting the values of the gauge boson mass $m_{Y}$. In panel (c), the correlation between $\text{Br}(\tau\to e\gamma)$ and the conversion rate $\mathrm{CR}(\mu^{-}\text{Al}\to e^{-}\text{Al})$ is presented, with a color gradient reflecting the values of the charged scaler mass $m_{\rho_{3}^{\pm}}$. This offers a direct indication of the dependence of the conversion rates on the mass of the new gauge boson in our model. The yellow shaded region in both panels marks the exclusion zone from the experimental limits set by the MEG II experiment [142].
@@
In Table 5, we present numerical cases for the heavy neutrino masses $m_{n_{4}}$ to $m_{n_{9}}$, also include the mass of the gauge boson $m_{Y}$ and the charged scalar mass $m_{\rho_{3}^{\pm}}$, both expressed in TeV, the corresponding branching ratios for processes $\mu\to e\gamma$, $\tau\to e\gamma$ and $\tau\to\mu\gamma$ and the estimated conversion rate $\mathrm{CR}(\mu^{-}\mathrm{Al}\to e^{-}\mathrm{Al})$. These values provide insight into the correlations between the LFV processes and the masses of the heavy neutrinos and gauge bosons within the context of our model.